OpenStudy (anonymous):

Please check: -3x^3-7x^2-3x-7 need to express as a product of linear factors and find all zeros X^2(-3x-7)(-3x-7) Zeros are -7/3

OpenStudy (solomonzelman):

this is an algebra 1 function, no other technics need to be applied. -3x^3-7x^2-3x-7=0 -3x^3-3x-7x^2-7=0 -3x(x^2-1)+(-7x)(x^2-1)=0 (-3x-7x)(x^2-1)=0 (-10x)(x^2-1)=0 -10x=0 x=0 x^2-1=0 x^2=1 x= -1, 1

OpenStudy (solomonzelman):

So the zeros are 0, 1, -1 the y intercept is just going to be the last constant, in this case -7. So zeros, 0, -1, 1 y-intercept, -7

OpenStudy (anonymous):

I don't quite follow how you took out the 3x and got -3x(x^2-1)+(-7x)(x^2-1)

OpenStudy (solomonzelman):

I factored it out, that's all.

OpenStudy (solomonzelman):

I mean not factored OUT, I just factored.

OpenStudy (anonymous):

Shouldn't I factor -3x^3-7x^2 as x^2(-3x-7) then have (-3x -7) as the other factor?

OpenStudy (solomonzelman):

you would get the same thing.

OpenStudy (anonymous):

How do I then get from (-3x-7) to the zeros of 0,-1,1?

OpenStudy (phi):

\[ -3x^3-7x^2-3x-7 \\ x^2(-3x -7) + (-3x-7) \\ (x^2 +1)(-3x-7) =0\]

OpenStudy (phi):

the term -3x -7=0 gives us -3x = 7 x = -7/3

OpenStudy (phi):

but the x^2 + 1 =0 gives us x^2 = -1 the square root of -1 is an imaginary number ... you get ± i (where "i" is short for sqr(-1) )

OpenStudy (phi):

If you are not studying complex roots, this might mean there is a typo in the original equation

OpenStudy (anonymous):

So I was on the right track with the -7/3 but didn't go far enough with the x^2+1?

OpenStudy (phi):

you should have gotten to this \[ x^2(-3x -7) + (-3x-7) \] to make it a bit more obvious, this is like factoring \[ x^2 A + A = (x^2 +1) A \]

OpenStudy (phi):

because minus signs are notorious for causing problems, we could at the very start \[ -3x^3-7x^2-3x-7= 0 \] multiply both sides by -1 (all terms!) to get \[ 3x^3+7x^2+3x+7=0 \] we will still get the same answer.... but it's less confusing (for me at any rate)

OpenStudy (anonymous):

I started out doing that then was confused by the reply. So. My zeros are -7/3, -1,1